OpenStudy (anonymous):

Jadwiga Skoron worked 9 hours at time- and- a- half pay. Her regular pay rate was $14.54 an hour. What was Jadwigs's total overtime pay?

OpenStudy (zpupster):

(9)(14.54*3/2)-----1 1/2=3/2 so then multiply by 9

OpenStudy (anonymous):

So $13.50 ?

OpenStudy (anonymous):

14.54/2=7.27 14.54+7.27=21.81 21.81*9=196.29

OpenStudy (anonymous):

Thank you